### v1

# Laminator: Verifiable ML Property Cards using Hardware-assisted Attestations

Abstract
Regulations increasingly call for various assurances from machine learning (ML) model providers about their training data, training process, and the behavior of resulting models during inference. For better transparency, companies (e.g., Huggingface and Google) have adopted model cards and datasheets which describe different properties of the training datasets and models. In the same vein, we introduce the notion of an inference card to describe the properties of a given inference (e.g., binding output to the model and its corresponding input). We collectively refer to these as ML property cards.
A malicious model provider can include false information in ML property cards, raising a need for verifiable ML property cards. We show how to realized them using property attestation, technical mechanisms by which a prover (e.g., a model provider) can attest different ML properties during training and inference to a verifier (e.g., an auditor). However, prior attestation mechanisms based purely on cryptography are often narrowly focused (lacking versatility) and inefficient. There is a need to efficiently attest different types properties across the ML model training and inference pipeline.
Recent developments make it possible to run and even train models inside hardware-assisted trusted execution environments (TEEs), which can provide highly efficient attestation. We propose Laminator, the first framework for verifiable ML property cards using hardware-assisted ML property attestations to efficiently furnish attestations for various ML properties for training and inference. It scales to multiple verifiers, and is independent of the model configuration.
I Introduction
Machine learning (ML) models are increasingly being deployed for various high-stakes applications like medical diagnosis, bank loans, and autonomous vehicles [9]. This has raised concerns about the different risks to data privacy, fairness and robustness [31]. Further, there are several emerging regulations to ensure that the training process and model’s behaviour during inference are as expected [22, 63, 14, 10, 12, 38, 32, 23]. For instance, a model provider may be required to convince a verifier (e.g., a regulator or potential customer) that a model meets the expected level of accuracy, fairness, privacy or robustness without disclosing the model [22, 23, 32, 38, 12] or demonstrate to an auditor that distributional properties of training data comply with regulations [14, 10] without disclosing the training data itself. The ability to demonstrate regulatory compliance without compromising confidentiality of models or training data will enable the development of a marketplace for ML models.
Several companies, including Google and Huggingface, have adopted model cards [43] and datasheets [18, 52] for better transparency of the properties of ML models or training datasets. We propse a similar notion, inference cards for describing the properties of a specific inference, e.g., binding the output of an inference, to the input and the model. Collectively, we refer to them as ML property cards which are short documents a containing information about ML properties such as intended use, risks and limitations, training procedure, datasets, and metrics for evaluation (e.g., accuracy, fairness, robustness) [43]. These ML property cards are generated by the model provider who trains and/or deploys the model. This requires trusting them to honestly provide the information in the ML property cards. While trusting reputable companies which open-source their models may be reasonable, we need to provide mechanisms to attest information in the ML property cards for regulatory compliance. Further, as seen on Huggingface, trained models and their corresponding property cards from potentially malicious entities are also available in model marketplaces [44]. Hence, there is a need for verifiable ML property cards to ensure that the information contained in them is neither false nor tampered with.
Huggingface provides a utility by which model providers can send their model to an endorser like Huggingface in order to compute some metrics, and create a verifiable property card that contains a signature to convince the verifier to ensure that a metric was verified by the endorser. However, this is not versatile, i.e., it is limited only to metrics, models and datasets implemented or allowed by the endorser. Hence, it is not possible to prove i) properties of sensitive datasets that cannot be disclosed to the endorser (e.g., demographic distribution, or certifying provenance of training dataset) ii) that a model was trained using a specific optimization metric (e.g., differential privacy [4], or group fairness [21]), or iii) that an output is the result of the model applied on a corresponding input during inference. We propose using ML property attestation, first introduced by Duddu et al. [11], which are technical mechanisms by which a prover (e.g., a model provider) can demonstrate relevant properties (including training-time and inference-time) of the model to a verifier (e.g., a client querying a model deployed as a service, a customer purchasing a model, or a regulatory authority).
Current ML property attestation schemes have various drawbacks in terms of accuracy [11], efficiency [11, 17], or scalability to multiple verifiers [11]. Further, prior works are limited to specific attestations such as proof of training (that a model was trained on some dataset) [17, 5, 57, 62], distributional property attestations (showing the training dataset of a model satisfies distributional properties) [11, 8], or showing that a model was trained with some optimization [56, 34, 50, 55, 68]. There is a need to efficiently attest various different types of properties during ML training or inference. We identify efficiency, scalability, and versatility as important criteria for attestation schemes.
Trusted Execution Environments (TEEs) already have the notion of hardware-assisted remote attestation to prove local system or software configuration to a remote verifier. The trusted computing research community has extended this to the notion of property attestation [54]. Recent developments by hardware vendors [26, 49], make it possible to run and train ML models efficiently inside TEEs. Hence, we identify TEEs as the choice for realizing efficient, scalable and versatile ML property attestation. We are the first to propose a software framework for hardware-assisted ML property attestation. We use TEEs to establish bindings between key properties of training dataset and the model during inference and training. These bindings are signed by the TEE’s secret attestation key, generating ML property attestations. When combined with signed external certificates from trusted entities, these attestations enable verifiers to draw conclusions about the model and training dataset properties during inference. These attestations can be then be used for verifiable ML property cards. We claim the following contributions: we
  1. 1.
    propose inference cards which include properties specific to each inference (e.g., binding input-model-output) and the collective of ML property cards to include model cards, datasheets, and inference cards (Section IV)
  2. 2.
    propose Laminator  [^1] We will make the code publicly available after publication., the first framework to generate verifiable ML property cards using hardware-assisted attestations throughout training and inference (Section V)
  3. 3.
    design, implement and evaluate four different types of attestations on Intel SGX as part of Laminator (Section VI)
II Background
We present the ML notations used in this work (Section II-A), details about model and dataset cards (Section II-B), and trusted execution environments (Section II-C).
II-A Machine Learning Notations
Consider an ML model architecture configuration $\mathcal{M}_{ar}$ indicating the number and types of layers, and a training configuration file $\mathcal{T}$ which includes the objective function, learning rate, weight decay, and other hyperparameters such as number of epochs. Additionally, consider a dataset $\mathcal{D}_{tr}\sim\mathbb{D}_{\textsc{p}}$ where $\mathbb{D}_{\textsc{p}}$ is an underlying distribution satisfying some distributional properties p (for example, the ratio of females in the distribution). Since, we use i.i.d sampling to get $\mathcal{D}_{tr}$, it satisfies p. We obtain model $\mathcal{M}$ by training as per the description in $\mathcal{T}$ using $\mathcal{M}_{ar}$ on $\mathcal{D}_{tr}$, i.e., $\mathcal{M}\leftarrow\text{Train}(\mathcal{D}_{tr},\mathcal{M}_{ar},\mathcal{T})$. During inference, $\mathcal{M}$ takes an input $\mathcal{I}$ and output $\mathcal{O}$, i.e., $\mathcal{O}$leftarrow $\mathcal{M}$($\mathcal{I}$). Additionally, $\mathcal{M}$ is evaluated with respect to accuracy, fairness or robustness on some unseen test dataset $\mathcal{D}_{te}$ sampled from same distribution as $\mathcal{D}_{tr}$. We assume that both $\mathcal{D}_{tr}$ and $\mathcal{D}_{te}$ are composed of data records $\mathcal{I}$, the corresponding classification label $y$, and some sensitive attribute $z$. We summarize different notations used in the rest of the paper in Table I.
TABLE I: Summary of notations and their descriptions.
Notation Description
$\mathcal{P}$ Prover
$\mathcal{V}$ Verifier
p Distributional property
$\mathcal{M}_{ar}$ Model architecture
$\mathcal{T}$ Training Configuration
$\mathcal{I}$ Input to a model
$y$ Classification label
$z$ Sensitive attribute
$\mathcal{O}$ Output from a model for $\mathcal{I}$
$\mathcal{M}$ Prover’s model to be attested
$\mathcal{D}^{prov}_{tr}$ Prover’s training dataset
$\mathcal{D}^{prov}_{te}$ Prover’s test dataset
$\mathcal{D}_{rob}$ Test dataset with adversarial examples
$\mathcal{A}_{acc}$ Test accuracy
$\mathcal{A}_{rob}$ Robust accuracy
$\mathcal{A}_{fair}$ Fairness metric
$\mathcal{A}^{ver}_{acc}$ Verifier’s measured accuracy
$\mathcal{A}^{ver}_{rob}$ Verifier’s measured robust accuracy
$\mathcal{A}^{ver}_{fair}$ Verifier’s measured fairness metric
$h(\cdot)$ cryptographic hash function
II-B Model Cards and Datasheets
To improve transparency into ML training and inference, companies have adopted model cards and datasheets. We describe them below.
Model cards are short documents containing additional information about different static model properties [43]. These include properties prior to deployment of the model as a service to clients to query via an API reference. They were designed to provide better transparency into the design of ML model which includes training and evaluation of the model prior to deployment for inference. These have been extensively adopted by different companies such as Google [20] and Huggingface [25]. Model cards contain a variety of additional information about ML properties: i) general propertiesincluding information about the model’s intended use, risks such as bias, and other limitations ii) training propertiesincluding training datasets used, training hyperparameters, and model architecture iii) evaluation propertiesincluding information about the different test datasets and metrics used for measuring accuracy, precision, recall, fairness, and robustness.
Datasheets describe different properties corresponding to the training dataset  [18, 52]. Datasheets, also known as dataset cards, have been used extensively by companies such as Huggingface [1]. They describe collection process and sources from which the raw data was collected, pre-processing including cleaning and labelling to obtain training dataset from raw data, intended use, composition of dataset, and demographic subpopulations and their distribution.
Model cards and datasheets are created by model providers who train the model and/or deploy them as a service. Hence, we have to trust the model providers for the information claimed in the cards.
II-C Trusted Execution Environments (TEEs)
TEEs [36] provide an environment in which security-critical code (trusted applications, or TAs) can run without outside interference from other TAs, or the rest of the system (the Rich Execution Environment, or REE, which contains the operating system and normal software). TEEs can be described in terms of three main properties: isolation, secure storage, and remote attestation [36]. Isolation refers to the inability for outside software to interfere with the operation of a TA, or to read its confidential state. Secure storage refers to persistent storage made available to TAs that resists interference by other software. Remote attestation refers to the ability of a TA to prove some aspect of a TEE’s state to a remote party, e.g., that the secret key of a particular keypair is held by a TA whose code has a particular cryptographic hash.
In this work, we will focus on the use of Intel’s Software Guard Extensions (SGX) [2] to enable TEEs, but the work equally applies to other TEE technologies that provide both isolation and remote attestation. SGX provides enclave-type TEEs, meaning that individual applications can instantiate a TEE (an enclave) in their own memory space, running a TA of their choice. SGX prevents the host application, the operating system, or any other code outside the enclave, from reading or writing to the enclave’s memory directly. Communication with the enclave is via well-defined entry and exit points, as well as by the TA reading and writing the host application’s memory outside of the enclave.
SGX enclaves can remotely attest themselves by obtaining a quote, which functions as a signature over TA-provided ‘user data’, along with a certification of the enclave identity that produced it. This enclave identity includes MRENCLAVE, a hash of the TA’s initial memory image, MRSIGNER, a hash of the public key that signed the TA, whether the enclave is in debug mode, versioning information, and various other data [2, EREPORT].
A verifier can evaluate the validity of the quote, and use its knowledge of the code whose hash is contained in the quote’s MRENCLAVE (and specifically, the circumstances under which it will produce a quote) to reliably determine properties of the TA’s state at the time that it produced the quote. In particular, if a particular TA produces quotes only over the result of a particular computation, then the verifier can be assured that such a computation was performed, and yielded the result contained in the quote. This can be used to prove arbitrary properties of data available to and computations performed by a TA.
II-D Gramine Library OS
Gramine [^2] https://gramineproject.io/ allows one to run unmodified applications inside an SGX enclave using a library OS, enabling the use of off-the-shelf software such as Python. This requires a manifest file which describes the application’s security configuration, including information on files that will be accessible, with or without integrity checking. By default, only files explicitly listed as trusted files (whose data is incorporated into the enclave hash at build time, and have their integrity checked on access), or allowed files (which can be accessed but are not subject to integrity checks) are accessible from within the enclave.
This occurs by the Gramine tooling automatically hashing the trusted files, and adding them to the enclave manifest file, which is incorporated into the SGX enclave’s MRENCLAVE. Thus, marking files as trusted prevents them from being silently replaced by a malicious host, as the subsequent hash verification will fail. This allows even security-critical files to be stored outside the enclave.
III Problem Statement
We propose the term ML property cards to collectively refer to declarations like model cards and datasheets used by model providers in the interest of transparency. Our goal is to design a framework to generate verifiable ML property cards with the help of ML property attestations for training and inference. This will help attest information in ML property cards (a) for regulatory compliance, and (b) is not false or tampered. We first present the adversary model (Section III-A), desiderata for an ideal attestation mechanism (Section III-B), and limitations of current attestation mechanisms (Section III-C).
III-A Adversary Model
We consider a prover ($\mathcal{P}$) that designs, trains or evaluates a model $\mathcal{M}$. During training, $\mathcal{P}$ wants to attest different properties of $\mathcal{M}$’s training to a verifier ($\mathcal{V}$). For instance, attesting that $\mathcal{M}$ was trained using $\mathcal{T}$, $\mathcal{M}_{ar}$, and $\mathcal{D}_{tr}$ (aka “proof of training”) which satisfies some required property $p$. After training, $\mathcal{P}$ evaluates $\mathcal{M}$ to check that is satisfies some minimal accuracy, fairness and robustness requirements prior to deployment. Once $\mathcal{P}$ deploys $\mathcal{M}$ as a service, clients can query the model by sending $\mathcal{I}$ via an API interface and obtain the corresponding $\mathcal{O}$. Here, $\mathcal{P}$ attests that $\mathcal{O}$ was generated $\mathcal{M}$ for a specific $\mathcal{I}$. In Section IV-A, we expand the notion of ML property cards to include a new category, inference cards to cover this type of inference-time attestation.
Since $\mathcal{P}$ is untrusted, they can try to fool the attestation by trying to attest themselves despite not meeting the requirements. For instance, making false claims despite not using the expected $\mathcal{T}$ or $\mathcal{M}_{ar}$, using a different dataset than $\mathcal{D}_{tr}$ to train $\mathcal{M}$ which does not satisfy $p$, incorrect values for accuracy, fairness, and robustness, or using a different $\mathcal{O}$ for inference. We assume that $\mathcal{P}$’s entire computing platform except for the TEE is untrusted, including the privileged code such as the operating system and hypervisor are controlled by $\mathcal{P}$.
We assume two roots of trust: (a) signature verification keys of external trusted organizations (e.g., CIFAR) which provides external certificates, e.g., for a test dataset, and (b) the manufacturer of the TEE (e.g., Intel) who certifies the keys used by enclaves to signi attestations. As long as $\mathcal{V}$ trusts these roots of trust, $\mathcal{V}$ can then derive trust in any attestation or external certificate that can be mapped to them.
III-B Desiderata for ML Property Attestations
We identify the following desiderata that an underlying ideal ML attestation scheme must satisfy:
  1. R1
    Efficient (low-overhead generation and verification of the attestation, even with large models and a wide range of useful ML properties).
  2. R2
    Scalable (supports large numbers of provers/verifiers).
  3. R3
    Versatile (able to prove a wide variety of claims, and new claims created with minimal effort.)
Our goal is to design a framework to generate verifiable ML property cards using an attestation mechanism which can satisfy all the above requirements.
[…]
Input: $i_{1},i_{2},\ldots$
Computation: $o_{1},o_{2},\ldots\overset{\$}{\leftarrow}f(i_{1},i_{2},\ldots)$
Output: $o_{1},o_{2},\ldots,\operatorname{Sign}_{\textrm{att}}(\texttt{'AttX'},g(i_{1},i_{2},\ldots,o_{1},o_{2},\ldots))$
Assertion: $P(i_{1},i_{2},\ldots,o_{1},o_{2},\ldots)$
Att 1 Attestation
where $a\overset{\$}{\leftarrow}\operatorname{Proc}()$ denotes probabilistic assignment (i.e. assigning to $a$ the value of a possibly-probabilistic procedure $\operatorname{Proc}()$).
[…]
III-C Limitations of Prior Work
A popular ML property attestation mechanism is using cryptographic primitives such as zero knowledge proofs (ZKPs) and secure multiparty computation (MPC). However, none of the prior attestation scheme using cryptographic primitives satisfy all the three requirements.
  1. 1.
    ZKPs are scalable and can support multiple verifiers. However, they are computationally expensive and do not efficiently extend to neural networks. For instance, proof of training using ZKPs for a simple logistic regression model on a dataset of $2^{18}$ data records with 1024 attributes, and a batch size of 2014, takes 72 seconds for one epoch including training and proof generation time [17]. Finally, ZKPs are lack versatility, and require properties that can be adapted to the ZKP scheme [17], such as complex nonlinear activation functions.
  2. 2.
    MPC requires online interactions between the prover and verifier, meaning that it lacks scalability and efficiency, and requires training the model for each verification [11].
IV Verifiable ML Property Cards
ML property cards currently cover properties about datasets in the form of datasheets, and about models in the form of model cards. However, they currently do not include properties during each inference after deployment. To this end, we propose the notion of inference cards (Section IV-A). We then present different ML property attestations for different verifiable ML property cards (Section IV-BIV-C, and IV-D). Finally, we discuss how $\mathcal{V}$ can combine different attestations to make meaningful conclusions (Section IV-E).
IV-A Inference cards
Inference cards include properties for each inference: i) attesting that an output was from a specific model and input ii) verifiable inference for integrity of the outputs using ZKPs for ML models [40, 24, 16] and large language models [61] iii) attesting that model explanations, vector indicating the influence of different attributes in input to generate output, corresponds to a specific model, input, and output iv) attesting that an processed information was from either a pre-processing or post-processing filter for a specific input.
Train Dataset ($\mathcal{D}_{tr}$)Data metrics(bias, size)Architecture($\mathcal{M}_{ar}$)Configuration($\mathcal{T}$)Trained ModelTest Dataset ($\mathcal{D}_{te}$)Accuracy, Fairness,RobustnessInputOutputDatasetAttestationProof of TrainingEvaluationAttestationInferenceAttestation
Figure 1: Overview of ML property attestations and their relation to different verifiable model cards: attestations for datasheet (orange), model card (blue), inference card (red), along with external certificates (cyan).
IV-B Attestations for Datasheets
We first present dataset attestations required for datasheets: distributional property attestation of $\mathcal{D}_{tr}$. We hereafter refer to a hash function as $h(\cdot)$. All the frequently used notations are in Table I.
Distributional Property Attestation attests that distributional properties p (such as gender ratio) of $\mathcal{D}_{tr}$ by providing an assertion that it has some value which can later be checked with some requirements by $\mathcal{V}$ [11]. This can either be done during training [11, 17] or later during inference [11] with Algorithm 3. We can have different distributional properties across: i) a single attribute (e.g., male-to-female ratio) ii) multiple attributes (e.g., male-to-female and white-to-non-white ratios) iii) joint properties (i.e., fraction of data records for a given classification label for different subgroups). For a given $\mathcal{D}^{prov}_{tr}$, Laminator specifies the properties to be attested and generates the following: p, h($\mathcal{D}^{prov}_{tr}$), $sign_{sk}$(h($\mathcal{D}^{prov}_{tr}$), p)). We present the attestation in Algorithm 2.
Input: $\mathcal{D}^{prov}_{tr}$
Output: p, $\operatorname{Sign}_{\textrm{att}}(\texttt{'DistAtt'},\textsc{p},\texttt{h}(\mathcal{D}^{prov}_{tr}))$
Assertion: $\textsc{p}(\mathcal{D}^{prov}_{tr})$
Attestation 2 Distributional Property Attestation (DistAtt)
IV-C Attestations for Model Cards
We now cover different attestations for model cards which includes: i) proof of training which includes attesting $\mathcal{T}$ and $\mathcal{M}_{ar}$ ii) evaluation attestations (accuracy, on $\mathcal{D}^{prov}_{te}$ iii) on $\mathcal{D}^{prov}_{te}$ iv) robustness on $\mathcal{D}^{prov}_{te}$ We indicate these in blue (Figure 1).
Proof of Training attests that $\mathcal{M}$ was trained on $\mathcal{D}^{prov}_{tr}$. This requires training $\mathcal{M}$ using a training configuration $\mathcal{T}$ on $\mathcal{D}^{prov}_{tr}$ inside the trusted hardware. The attestation in shown in Algorithm 3.
Input: $\mathcal{D}^{prov}_{tr}$, $\mathcal{T}$
[…]
where $\mathcal{M}\leftarrow\operatorname{Train}(\mathcal{D}^{prov}_{tr},\mathcal{T},\mathcal{M}_{ar})$
Attestation 3 Proof of Training (PoT)
Accuracy attestation indicates that $\mathcal{M}$ has a certain accuracy $\mathcal{A}_{acc}$ as measured using some dataset $\mathcal{D}^{prov}_{te}$. This ensures i) there are no false claims about $\mathcal{M}$’s $\mathcal{A}_{acc}$ ii) $\mathcal{M}$was evaluated on (a standardized) $\mathcal{D}^{prov}_{te}$ with no tampering. Laminator, in addition to $\mathcal{A}_{acc}$, generates $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$) along with a signature binding them, i.e., $Sign_{\textrm{att}}$($\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{acc}$). We use an indicator function $\mathbb{I}_{\{a=b\}}$ that takes the value $1$ if $a=b$, and 00 otherwise. We summarize this in Algorithm 4.
Input: $\mathcal{M}$, $\mathcal{D}^{prov}_{te}$
Output: $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{acc}$, $\operatorname{Sign}_{\textrm{att}}(\texttt{'AccAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}^{prov}_{te}),\mathcal{A}_{acc})$
Assertion: $\mathcal{A}_{acc}=\frac{1}{N}\sum_{(x,y)\in\mathcal{D}^{prov}_{te}}^{N}\mathbb{I}_{\{\mathcal{M}(x)==y)\}}$
Attestation 4 Accuracy attestation (AccAtt)
Fairness Attestation extends accuracy attestation to evaluate fairness of $\mathcal{M}$ on $\mathcal{D}^{prov}_{te}$ using different metrics such as accuracy parity, false positive and negative rates across subgroups. We refer to the fairness metric being evaluated as $\mathcal{A}_{fair}$. We attest that $\mathcal{M}$ satisfies $\mathcal{A}_{fair}$ on $\mathcal{D}^{prov}_{te}$ in Algorithm 5. We only indicate demographic parity but any other fairness metric can be used in its place.
Input: $\mathcal{M}$, $\mathcal{D}^{prov}_{te}$
Output: $\mathcal{M}$), h($\mathcal{D}^{prov}_{te}$), $\mathcal{A}_{fair}$,$\operatorname{Sign}_{\textrm{att}}(\texttt{'FairAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}^{prov}_{te}),\mathcal{A}_{fair})$
Assertion: $\mathcal{A}_{fair}=|P(\mathcal{M}(X)=0|S=0)-P(\mathcal{M}(X)=0|S=1)|$ where $(X,Y,S)$ $\in\mathcal{D}^{prov}_{te}$
Attestation 5 Fairness Attestation (FairAtt)
Robustness Attestation attests that $\mathcal{M}$ satisfies some acceptable level of robustness indicated by robust accuracy ($\mathcal{A}_{rob}$) on a test dataset containing adversarial examples ($\mathcal{D}_{rob}$). This proceeds in two steps: first, we generate $\mathcal{D}_{rob}$ along with its hash (h($\mathcal{D}_{rob}$)) inside the enclave. Since $\mathcal{D}_{rob}$ is generated locally, its hash acts as the certificate during verification. This certified that $\mathcal{D}_{rob}$ was created by adding adversarial examples generates with a perturbation of $\epsilon$ to $\mathcal{D}^{prov}_{te}$. In the second step, we use accuracy attestation (Algorithm 4) but using $\mathcal{D}_{rob}$ instead of $\mathcal{D}^{prov}_{te}$, yielding our robustness output.
Generating and Certifying $\mathcal{D}_{rob}$:
Input: $\mathcal{D}^{prov}_{te}$, $\mathcal{A}_{rob}$, $\epsilon$
Output: $\mathcal{D}_{rob}$, h($\mathcal{D}_{rob}$), h($\epsilon$), h($\mathcal{A}_{rob}$), $\operatorname{Sign}_{\textrm{att}}(\texttt{'RobustAttA'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}_{rob}),\mathcal{A}_{rob})$
Attesting Robustness:
Input: $\mathcal{M}$, $\mathcal{D}^{prov}_{te}$
Output: $\mathcal{M}$), h($\mathcal{D}_{rob}$), $\mathcal{A}_{rob}$, $\operatorname{Sign}_{\textrm{att}}(\texttt{'RobustAttB'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{D}_{rob}),\mathcal{A}_{rob})$
Assertion: $\mathcal{A}_{rob}=\frac{1}{N}\sum_{(x,y)\in\mathcal{D}_{rob}}^{N}\mathbb{I}_{\{\mathcal{M}(x)==y)\}}$
Attestation 6 Robustness Attestation (RobustAtt)
IV-D Attestations for Inference Cards
We attest the correctness of input-model-output triples, which can be used to produce inference cards.
Input-Model-Output Attestation (IOAtt), shown in Algorithm 7 attests that the output $\mathcal{O}$ was generated from $\mathcal{M}$ for a given input $\mathcal{I}$. This ensures i) integrity against malicious $\mathcal{P}$ who can tamper with $\mathcal{O}$ ii) non-repudiation to prevent a malicious client from claiming that $\mathcal{O}$ does not belong to $\mathcal{I}$ iii) false claim that $\mathcal{O}^{\prime}$ is from $\mathcal{M}$ for $\mathcal{I}$ instead of $\mathcal{O}$.
Input: $\mathcal{I}$, $\mathcal{M}$
Output: $\mathcal{O}$, $\mathcal{M}$), $\mathcal{I}$), h($\mathcal{O}$), $\operatorname{Sign}_{\textrm{att}}(\texttt{'IOAtt'},\texttt{h}(\mathcal{M}),\texttt{h}(\mathcal{I}),\texttt{h}(\mathcal{O}))$
Assertion: $\mathcal{O}=\mathcal{M}(\mathcal{I})$
Attestation 7 Input-Model-Output Attestation (IOAtt)
IV-E Chain of Attestations and External Certificates
In addition to the above attestations, there could be additional certificates which allow verifiers to check the integrity of datasets and trained models. We consider two different external certificates: i) training dataset ii) test dataset, which certify the quality of these dataset as indicated by some trusted organization (e.g., CIFAR). Combining these external certificates can help $\mathcal{V}$ conclude that the model was trained and evaluated on dataset considered good for a particular task.
The assertions of different attestations and external certificates by themselves might not be useful in practice. For instance, input-model-output attestation suggests that for an input $\mathcal{I}$, $\mathcal{M}$ generates the output $\mathcal{O}$. However, this information is useful only if $\mathcal{V}$ knows how good $\mathcal{M}$ is in terms of accuracy, fairness or robustness. In order to use attestations in practice, we chain together different attestations included to obtain verifiable ML property cards that allow us to arrive at a wider range of conclusions.
As a motivating example, consider a company which has outsourced training to a third party before deploying some model ($\mathcal{M}$) as a service to clients. Using input-model-output attestation, the company can indicate that for their input $\mathcal{I}$, $\mathcal{M}$ generated $\mathcal{O}$ (Algorithm 7). Additionally, they have to attest that $\mathcal{M}$ is a good model in terms of having “good” accuracy on on some $\mathcal{D}_{te}$ using accuracy attestation (Algorithm 4). This helps the clients or $\mathcal{V}$ conclude that their $\mathcal{O}$ is indeed from a good model. Additionally, including other attestations such as fairness (Algorithm 5) and robustness can help show that $\mathcal{M}$ satisfies other properties as required by different applications. In all these cases, including the certification for $\mathcal{D}_{te}$ will also help conclude that the metrics are computed on a standardised test dataset which is certified by a trusted organization.
Combining several training-time attestations (e.g. Algorithms 3 and 2) can help to arrive at the following conclusions: $\mathcal{M}$ was trained on $\mathcal{D}_{tr}$ satisfying distributional properties p. However, we can combine training-time attestation with inference-time attestations to also reach conclusions about the properties of models after training (accuracy, fairness and robustness). Hence, in addition to training attestation, we can additionally prove that $\mathcal{O}$ was generated from $\mathcal{M}$ for $\mathcal{I}$, where $\mathcal{M}$ was trained on was trained on $\mathcal{D}_{tr}$ satisfying p, and $\mathcal{M}$ satisfies the required {accuracy, fairness, robustness}.
V Laminator: Framework Design
We present Laminator, a framework to generate verifiable ML property cards using ML property attestation. We identify hardware-assisted ML property attestations using Trusted Execution Environments (TEEs) as the choice for realizing efficient, scalable and versatile ML property attestation. TEEs already have the notion of hardware-assisted remote attestation to prove local system or software configuration to a remote verifier. The trusted computing research community has extended this to the notion of property attestation [54], and recent developments by hardware vendors [26, 49], make it possible to run and train ML models efficiently inside TEEs. We are the first to propose a software framework for hardware-assisted ML property attestation.
Laminator establishes bindings between key components: the model and its inputs/outputs, the model and its accuracy with respect to a test dataset, the model and its training dataset and configuration, and distributional properties of a dataset. Bindings are signed with the TEE’s secret attestation key, yielding attestations. As these can be generated by anyone with the requisite hardware, and validated by anyone, this approach is inherently scalable. They enable verifiers to draw conclusions about the model and training dataset properties during training and inference. We implement the different attestations described in Section IV. We revisit the completeness and how to extend Laminator for new attestations in Section VIII. We now discuss the design and implementation of Laminator for $\mathcal{P}$ and $\mathcal{V}$.
SGX EnclavesMeasurerPytorchPythonGramine LibOSAttestationsEndorserEndorsementsAssertionBundleIntel APICertificatesVerifierDCAPProperty Cards
Figure 2: Overview of Laminator’s design. Components already existing or part of the infrastructure are indicated in gray while components implemented are indicated in orange. There can be different enclaves for different attestations which are generated by changing the “measurer” (in green). Dashed lines correspond to components for endorser while solid lines correspond to components from TEE.
We present an overview of Laminator’s implementation in Figure 2. An SGX enclave runs a Python script (“measurer”) that measures properties of a model, dataset, or inference, and produces an attestation of a property card fragment (i.e. a chunk of JSON with the same structure as model card metadata that will be combined with other such fragments to produce the full property card). A combination of external certificates and various enclaves’ attestations is referred to as the assertion bundle. Since it is possible to endorse new enclaves without updating all verifiers’ software, this approach is inherently flexible. Furthermore, provisioning certificates are included in the enclave platforms using an Intel API to later check for validity of the attestations. We also consider assertions by endorsers (e.g., CIFAR, Huggingface) of the existing kind, who can be trusted to provide external certificates asserting that some publicly available models or datasets are of good quality, or simply to name a dataset. During verification, the assertion bundle can be passed to $\mathcal{V}$ who verifies the attestations using certificates from the TEE manufacturer (in this case, obtained from an API provide by Intel and used by the DCAP Quote Verification Library), and combines their assertions with those of the endorsers to obtain a set of verified ML property cards.
V-A The Prover
We implement Laminator’s prover with the aid of the Gramine framework [37], which allows the execution of unmodified applications inside an SGX enclave. This allows Laminator to produce attestations of properties computed using off-the-shelf libraries such as PyTorch [51].
We implement the attestations from Section IV. The different inputs and outputs from different attestation enclaves are shown in Figure 3. The dataset enclave takes $\mathcal{D}_{tr}$ as an input and generates the distribution along with DistAtt. The training enclave takes the $\mathcal{D}_{tr}$, $\mathcal{M}_{ar}$, and $\mathcal{T}$ as input, to train a model inside the enclave, and output the trained model $\mathcal{M}$ and PoT. Given $\mathcal{M}$ and $\mathcal{D}_{te}$, the metric enclave can generate different metrics for accuracy, fairness, and robustness, along with their corresponding attestations as an output. These metrics are used to evaluate $\mathcal{M}$ before deployment. Finally, once $\mathcal{M}$ is deployed, it can be run inside an inference enclave such that for each input, we get the corresponding $\mathcal{M}$’s output along with IOAtt.
SGX EnclavesTraining Dataset($\mathcal{D}^{prov}_{tr}$)Training and ModelConfiguration ($\mathcal{T}$+$\mathcal{M}_{ar}$)Dataset Enclave….Training Enclave….DistributionDistAttModelPoTInputInference Enclave….Metric Enclave….Test Dataset($\mathcal{D}_{te}$)MetricAccAttFairAttRobAttOutputIOAtt
Figure 3: Inputs and outputs for different attestation enclaves. Laminator includes four different enclaves which are indicated in orange. Attestations are in ellipse, outputs are in blue, and inputs to enclaves are white. We indicate measurer in green.
The attestations are generated from a custom ML library built atop PyTorch. A hash of the Python code responsible for these attestations, along with PyTorch and all other files used by the enclave, is embedded into the enclave using the Gramine framework, which provides a library OS that allows the integrity of the Python code to be verified when read from disk, incorporating it into the MRENCLAVE value included in SGX attestations.
Laminator includes a wrapper for Python’s open library call, which is used to read input files into memory, hashing them at the same time, allowing PyTorch to randomly access measured files without risk of time-of-check-time-of-use (TOCTOU) vulnerabilities. We implement an additional mechanism to check the resulting hash against a manifest, allowing for a single short measurement of large directories of input files without the need to keep them all in memory.
After the computation is complete, a hash of the output is incorporated into the user data field of an SGX report, which is then used to obtain a DCAP quote [3] that can be verified by remote parties. This output takes the form of a JSON string representing a fragment of the model card metadata [^3] https://github.com/huggingface/hub-docs/blob/main/modelcard.md, where a model is named with the hash of its file.
V-B The Verifier
Laminator’s verifier takes two main inputs:
  • A set of serialized attestations in JSON format, each containing an SGX quote—representing the attestation signature from Section IV—and base64-encoded output.
  • A set of certifications relating SGX enclave identities to the property that they attest; in our case, a hard-coded mapping from SGX enclave measurements to JSON templates for the property card metadata that they are allowed to assert.
The attestations are designed in such a way that they directly map to YAML-formatted property card metadata, similar to that included in existing Model Cards. This greatly simplifies the verifier, which needs only to check that the attested JSON corresponds to the properties included in the corresponding certification.
$\mathcal{V}$ validates each quote, validating it against the output of the respective attestation, then using the enclave measurement from the quote to look up the respective certification, which contains a JSON structure that serves as a template for the payload, and which is checked according the following rules:
  • null in the certification matches anything in the payload
  • Dictionaries in the certification match dictionaries (or arrays of dictionaries) in the payload whose keys match those of the certification, and whose values match the values of the certification
  • Strings, booleans, and integers in the certification match identical values (or arrays of identical values) in the payload
  • Other values are not allowed in the certification
Specifically, we do not allow the use of floating-point values in order to avoid rounding issues due to verification; these must be represented as JSON strings.
For each model referenced by an attestation, $\mathcal{V}$ produces a model card, formed by taking the union of the various model card metadata fragments to produce a single JSON dictionary that is reformatted into the YAML contents of the model card model-index field, which contains experimental data in a structured form.
V-C Authorization
Enclaves must only generate attestations of properties that they are not certified for and $\mathcal{V}$ should not trust an unknown enclave or trust an attestation from an enclave which it is not certified to make. To check for this, an endorser can provide an external certificate for an enclave, in addition to those for datasets or models, to indicate which properties each enclave is allowed to generate. These external certificates can be included in the assertion bundle and validated during the verification, though in our proof-of-concept they are hardcoded into the verifier.
VI Evaluation
We present the experimental setup which includes datasets, model architecture, training configuration, and metrics (Section VI-A). We then present a performance evaluation of different attestations (Section VI-B).
VI-A Experimental Setup
We now present different datasets, model architecture and training configuration for each of the dataset, and metrics for evaluation.
Datasets. We consider two datasets, CIFAR and CENSUS, for evaluating different ML property attestations which will help design verifiable ML property cards. Both the datasets are standard benchmark datasets used in most of the ML literature. We describe these datasets below.
CIFAR consists of 60000 images of 32 $\times$ 32 colour images belonging to one of ten classes of different objects: airplane, bird, cat, deer, automobile, dog, frog, horse, ship, and truck. The dataset contains 6000 images per class. We use 50000 training images and 10000 test images.
CENSUS consists of 48842 data records with 42 attributes about individuals from the 1994 US Census data obtained from UCI Machine Learning dataset repository. It includes sensitive attributes such as sex and race of the participant. Other attributes include marital status, education, occupation, job hours per week among others. The classification task is to estimate whether the individual’s annual income is at least 50,000 USD. We used 24,000 training data records and 24,000 testing data records. This dataset has two sensitive attributes, race and sex, which we use for attesting distributional properties and fairness.
Model Architecture. We now describe the model architecture used for both the datasets. For CIFAR, we use a convolutional neural network (CNN) model with three convolution layers of dimensions: [32, 64, 128]. Each being a filter with padding of 1 and kernel size of 3x3. After each convolution layer, there is a maxpool layer of size of 2x2 and a ReLU activation function and a batch normalization layer. The final layer has 512 nodes with a dropout layer with 0.25 probability of dropout. For CENSUS, we use a multi-layered perceptron (MLP) with three layers with each hidden layer of dimensions: [32, 64, 32]. Further, we use Tanh activation function.
Training Configuration. We train the model on CIFAR with 52 epochs with a learning rate of 1e-3, using Adam optimizer, and a batch size of 64 achieving accuracy of 81.31%. For CENSUS, we train the model with 10 epochs, learning rate of 1e-3, Adam optimizer, and batch size of 256 achieving accuracy 84.43%. We chose the number of epochs based on when the accuracy converges.
Metrics. We evaluate Laminator, breaking down the computation into four phases:
  • Startup which includes time spent starting the Python runtime, and in the case of Laminator, initializing the Gramine-based SGX enclave. This is a one-time-per-enclave-instance cost that is amortizable over multiple attestation requests.
  • Data pre-processing which includes time spent reading the data from disk, as well as performing any measurements or preprocessing.
  • Computation which includes the time spent on the inference, training, or property measurement operation.
  • Attestation which includes the time spent performing the attestation.
VI-B Performance Evaluation
We present the performance numbers for different attestations across the three ML property cards.
Attestations for Datasheets. We measure the performance of two attestations: distribution attestation, and proof of training.
Distribution Attestation. We evaluate distributional properties only for CENSUS dataset since it has sensitive attributes, race and sex. We generate attestations indicating i) the distribution of a sensitive attribute $z$ and ii) the distribution of $z$ given the classification label $y$ (denoted $z|y$ in Table II) We present the performance results for distributional properties of both race and sex attributes in Table II.
TABLE II: Performance of distributional property attestation.
Operation CENSUS-Race CENSUS-Sex
$z$ $z|y$ $z$ $z|y$
Startup 80 $\pm$ 9 s 80 $\pm$ 7 s 93 $\pm$ 11 s 75 $\pm$ 0.44 s
Preprocessing 330 $\pm$ 23 ms 320 $\pm$ 3.9 ms 370 $\pm$ 4.2 ms 320 $\pm$ 1.7 ms
Computation 14 $\pm$ 0.5 ms 27 $\pm$ 0.16 ms 15 $\pm$ 0.11 ms 27 $\pm$ 0.16 ms
Attestation 35 $\pm$ 71 ms 5.9 $\pm$ 0.15 ms 6.5 $\pm$ 0.55 ms 5.8 $\pm$ 0.079 ms
We find that the time for computation and attestation are low indicating good efficiency regardless of whether the sensitive attribute is race or sex. This is observed as the lack of significant difference in the time required to compute $z$ and $z|y$ of CENSUS-Race and CENSUS-Sex. Further, as expected, the computation time under the column for $z$ is less than $z|y$, due to the additional conditioning on the true classification labels $y$.
Attestations for Model Cards. We measure the performance of two attestations for verifiable model cards: proof of training, and accuracy attestations. We leave the evaluation of robustness and fairness attestations as they are similar to accuracy attestation but using different metrics.
Proof of Training. We train models for CIFAR and CENSUS inside Intel SGX and measure the time taken. We present the time averaged over ten runs indicated in milliseconds in Table III.
TABLE III: Performance of proof of training.
Operation CIFAR CENSUS
Startup 81 $\pm$ 0.82 s 71 $\pm$ 0.2 s
Preprocessing 4400 $\pm$ 24 ms 1300 $\pm$ 7 ms
Computation 3060 $\pm$ 120 s 36 $\pm$ 3.9 s
Attestation 5.4 $\pm$ 0.15 ms 5.8 $\pm$ 0.13 ms
We show that training an entire model for CIFAR takes $\sim$51 minutes while CENSUS takes $\sim$6 minutes. This is significantly more efficient than proof of training using ZKPs which is limited to simple models and datasets [17].
Accuracy Attestation. We measure the performance for both CIFAR and CENSUS datasets with 10,000 and 24,000 data records respectively. We present the results in Table IV.
TABLE IV: Performance of accuracy attestation.
Operation CIFAR (ms) CENSUS (ms)
Startup 36 $\pm$ 0.17 s 71 $\pm$ 0.17 s
Preprocessing 290 $\pm$ 2 ms 1300 $\pm$ 8 ms
Computation 3500 $\pm$ 100 ms 120 $\pm$ 3 ms
Attestation 5.7 $\pm$ 0.2 ms 5.7 $\pm$ 0.02 ms
The results show that the computation time for CIFAR is higher than CENSUS which is expected given the size of each CIFAR image is 32$\times$32$\times$3 compared to tabular data record with 42 attributes. Despite this, it takes only 3.5s to compute test accuracy on CIFAR while 0.1s for CENSUS.
Attestations for Inference Cards. We measure the performance of input-model-output attestation for verifiable inference cards.
Input-Model-Output Attestation. We evaluate the performance of input-model-output attestation by averaging 10 runs of 100 inferences. We present these results in Table V. For 100 inferences, the computation only takes 71ms for CIFAR10 compared to 21ms for CENSUS.
TABLE V: Performance of input-model-output attestation.
Operation CIFAR CENSUS
Startup 32 $\pm$ 0.19 s 71 $\pm$ 0.19 s
Preprocessing 0.50 $\pm$ 1.7 ms 6.3 $\pm$ 0.10 ms
Computation 71 $\pm$ 41 ms 21 $\pm$ 0.40 ms
Attestation 6.6 $\pm$ 1.1 ms 5.9 $\pm$ 0.20 ms
Summary. These findings highlight the efficiency of Laminator, with minimal overhead incurred during attestation generation. Notably, this initial startup cost is followed by the ability for limitless verification by any number of parties, affirming the scalability of our approach. Furthermore, TEEs exhibit versatility in training various models, exemplified by their successful implementation with complex architectures like AlexNet. In contrast, zero-knowledge proofs, constrained by model-dependent proofs, are restricted to simpler models.
VI-C Revisiting Requirements
We now indicate how the Laminator satisfies different requirements from Section III-B.
Efficiency. Running models inside SGX to generate attestations incurs a low overhead as, instead of developing custom ML pipeline, we can use high performance libraries like Pytorch along with the Intel’s AMX extensions. As shown above, $\mathcal{P}$ has some initial overhead but once started, multiple attestations can be generated with minimal overhead.
Once the attestation is made, the cost of verification is low, involving only certificate checking, and does not depend on the size of the model. Hence, Laminator satisfies R1.
Scalability. For each prover, attestation is done once and then can be verified independently by multiple verifiers. This process is non-interactive, and so the complexity of generating and verifying the quote does not depend on either the size of the model or the number of verifiers. Finally, the assertion bundle can combine many different attestations from different enclaves, which are then converted into a standard format; this means that Laminator can scale to large numbers of independent provers. All these design choices result in Laminator satisfying R2.
Versatility. As Laminator’s properties are measured for attestation purely in software, while hardware being used only to isolate this software, it is easy to extend Laminator to new properties. A developer can write a new Python script measuring the desired property. The enclave is built and its measurement certified, e.g., by a trusted endorser, or by hardcoding it into the verifier by its developer, enabling $\mathcal{V}$ to ensure that the new enclave can be trusted to attest its claimed assertions. This is aided by the attestation and certification formats, which can be directly mapped to YAML-formatted property card metadata, meaning that $\mathcal{V}$ need not be modified to introduce new metrics or ML property cards. Hence, Laminator satisfies R3.
VII Related Work
ML for Auditing. Several ML based techniques can be used for auditing ML models. For instance, membership inference attacks have been repurposed to be used for auditing by checking if data, requested to be removed as part of right to be forgotten, belongs to training data [58, 42]. However, such membership inference attacks have been shown to be unreliable and thereby lack effectiveness [53, 35]. Similarly, distribution inference attacks can be repurposed to be used for inference-based property attestation [11, 30]. However, these approaches lack effectiveness as well [11]. Further, Jia et al. [28] propose proof of training to identify whether a model was trained on a specific training dataset. However, several subsequent works have shown that these approaches can be easily evaded and are not effective [15, 69]. Hence, current ML based approaches for auditing are not reliable and robust.
Certified ML. Several prior works provide a “certificate” which is basically a bound on the accuracy in presence of adversarial examples (see a survey by Li et al. [39]). Further, implementation of differential privacy could have bugs due to which the actual privacy bounds are not the same as the claimed bounds. Several works have been proposed to audit differentially private models [60, 27, 48]. Finally, bounds to guarantee group and individual fairness can also be generated [29, 33, 65]. These bounds on different ML properties can be generated using randomized smoothing over the inputs or model parameters, or using formal theory. However, the threat model for these approaches are different from ours: these works implicitly assume that the prover who generates the proofs is honest while we assume the prover can be malicious. We instead rely on hardware as the root of trust.
Cryptographic Primitives for Auditing. ZKPs have been extensively used for verifiable inference to ensure that the model’s output has not been tampered [40, 16, 67, 24, 70, 19]. ZKPs for training are still in early stages and have been used for proof of training for different models [62, 57, 17, 5, 13]. However, this is not very efficient as proofs have to be generated for each of the training epoch which limits it to simple ML models [57, 17]. Extending these approaches to neural networks is an open problem as it incurs a significant overhead [62, 5]. Similarly, MPC for proof of training requires to train model for each attestation thereby incuring a high overhead [11]. Further, cryptographic primitives can verify properties such as fairness [57, 34, 50, 55, 68] and privacy [56]. Srivastava et al. [59] propose a verifiable training scheme aims to ensure determinism, mitigating hardware discrepancies and improving performance for auditing [17]. Finally, MPC has been used to verify entire training and inference pipeline but require multiple interactions [41].
Confidential Computing for ML. TEEs have been extensively explored for both training [45] and inference of ML models [46] (see survey by Mo et al. [47]). Prior works have relied on different optimizations including model partitioning to get ML models to execute within TEEs. Further, TEEs can be combined with GPUs for effiicent execution by using additional cryptographic operations to check for the integrity of the data transferred between TEE and GPU [64]. However, most these worst focus on ensuring confidentiality of the model and inputs during inference or training datasets. None of them use TEEs for integrity to furnish attestations.
VIII Discussion and Summary
Completeness. We consider three types of ML property cards and different attestations for each of them as illustrative examples in Section IV. However, we do not claim them to be complete. There could be other attestations and ML property cards but Laminator can be extended to account for them (discussed in Section V-A). For instance, ecosystem cards which considers a combination of multiple models trained on different datasets [6]. Hence, they comprise of multiple datasheets and model cards. However, Laminator can make them verifiable using verifiable model cards and datasheets.
Attacks on Intel SGX. Several prior works have shown that Intel SGX is susceptible to a wide range of side channel attacks by which an adversary can violate confidentiality [66]. However, we assume that model provider already has access to the training dataset and model details. There is no requirement for confidentiality but they use Laminator to show that their training and inference process is as expected. Hence, leakage via side channels is not a concern in our setting. The exception to this is that leakage of attestation keys will allow violation of the integrity of the attestation [7, 66]. Hence, Laminator relies on the ability of the platform developer to ensure the integrity of the attestation, using hardware and software mechanisms to prevent subversion e.g. by using side-channel or other attacks to extract attestation keys. However, this is an orthogonal problem to our own and the subject of active research and development effort on the part of TEE manufacturers.
Improving Efficiency with GPUs. We chose TEEs as the new variants have larger memory and rely on AMX extensions which can allow for faster inference and training. To improve efficiency, similar to prior works [64, 71], we can outsource the execution of some of the model layers to a GPU. While these works focus on confidentiality, we have to check for integrity of the computation done on the GPU. Further, using the improvements in hardware to combine GPUs with TEE (e.g., Nvidia’s H100 has a new TEE with a GPU), $\mathcal{P}$ can adapt Laminator to the new hardware to run both training and inference securely inside TEEs to improve efficiency [49]. This can allow to evaluate larger models including large language models.
Summary. ML property cards are widely adapted by companies such as Huggingface. There is a need to ensure the information claimed in these cards are correct creating the need for verifiable ML property cards. We propose Laminator, the first framework to generate verifiable ML property cards using hardware-assisted ML property attestations. We evaluate Laminator and show that it satisfies efficiency, scalability and versatility.
IX Acknowledgements
This work is supported in part by the Research Council of Finland (decision 339514), Intel Labs, and the Government of Ontario. Views expressed in the paper are those of the authors and do not reflect the position of the funding agencies.